# Towards Robust Infrared Small Target Detection: A Feature-Enhanced and Sensitivity-Tunable Framework

### v2

@@ 5 Conclusion
This manuscript proposes a feature-enhanced and sensitivity-tunable (FEST) framework, aiming to provide an efficient and universal optimization framework for infrared small target detection tasks. Specifically, we propose an edge enhancement difficulty mining (EEDM) loss, which guides the network to focus more on challenging regions and edge features during training, thereby promoting fine-grained learning and positive optimization of the model. Meanwhile, we adopt a multi-scale fusion strategy, which enhances the network’s perception and adaptability to multi-scale features of multi-size targets by integrating feature information from different scales. In addition, to fully utilize the probability mask information, we design an adjustable sensitivity (AS) strategy for network post-processing. This strategy employs a tunable dual threshold mechanism, which significantly improves the detection rate of infrared small targets while maintaining the segmentation accuracy of the model. Extensive experimental results demonstrate that the proposed FEST framework achieves significant performance gains and exhibits strong robustness on multiple public datasets. Notably, the lightweight LW-FEST framework achieves notable performance gains with minimal additional resource consumption.

### v3

@@ 5 Conclusion
This manuscript proposes a feature-enhanced and sensitivity-tunable (FEST) framework, aiming to provide an efficient and universal optimization framework for SIRST detection tasks. Specifically, we propose an edge enhancement difficulty mining (EEDM) loss, which guides the network to focus more on challenging regions and edge features, thereby promoting fine-grained learning and positive optimization of the model. Meanwhile, we adopt a multi-scale fusion strategy, which enhances the model’s perception to multi-scale features of multi-size targets. In addition, to fully utilize the probability mask information, we design an adjustable sensitivity (AS) strategy for network post-processing. This strategy employs a tunable dual threshold mechanism, which significantly improves the detection rate of infrared small targets while maintaining the segmentation accuracy of the model. Extensive experimental results demonstrate that the proposed FEST framework achieves significant performance gains and exhibits strong robustness on multiple public datasets. Notably, the lightweight LW-FEST framework achieves notable performance gains with minimal additional resource consumption.
In this study, although we provide a lightweight LW-FEST, the higher-accuracy FEST still inevitably brings certain computational costs because it employs a multi-scale fusion strategy. In future work, we will develop an adaptive detector in combination with a difficulty prediction mechanism. The detector will maintain lightweight inference on easy samples and dynamically activate more complex modules for hard samples, thereby reducing the overall inference burden. In addition, we plan to extend the current single-frame detection framework to multi-frame detection tasks to fully exploit temporal information and further improve detection performance. Notably, in practical scenarios, small targets often exhibit gradually emerging dynamics. Therefore, we will explore incorporating LSTM and other temporal prediction techniques [sultan2025intelligent, raja2025design, raja2025design2] to combine detection with forecasting of the target’s future position and shape, thereby providing decision support for early warning, interception, and observation-resource scheduling.